What is the equation of a line that has a slope of -3 and passed through (1, -2)

Answer by fractalier(6550)

One way to do this is to merely plug this info into
y = mx + b
to determine b...we get
-2 = -3(1) + b
-2 = -3 + b
b = 1
so that our line is
y = -3x + 1